# High-Throughput Low-Cost Segmentation of Brightfield Microscopy Live Cell Images

### v1

@@ Abstract
Notably, the model was trained primarily on bright-field images with limited exposure to phase-contrast microscopy (<<10%), yet it generalized effectively to the phase-contrast LIVECell dataset, demonstrating modality, robustness and strong performance. This highlights its potential for real-world laboratory deployment across imaging conditions.
The model requires minimal compute power and is adaptable using basic deep learning setups such as Google Colab, making it practical for training on other cell variants. Our pipeline outperforms existing methods in robustness and precision for bright-field microscopy segmentation. The code and dataset are available for reproducibility  [^1] The full dataset, code and the complete set of survey questions used in this study may be provided upon reasonable request to the corresponding author.
@@ 1 Introduction
Importantly, while our training set was composed primarily of bright-field images, it included a small proportion $(<10\%)$ of phase-contrast images, allowing us to evaluate the model’s robustness across imaging modalities. The model generalized effectively to phase-contrast images in the LIVECell dataset, demonstrating its adaptability and cross-modality potential.
@@ 2 Related Work
The work by [37] represents a significant shift toward annotation-free segmentation through optical flow-based pseudo-labeling. While achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images, our evaluation revealed critical limitations for bright-field applications: (1) processing times of 50–60 seconds per image due to iterative optical flow computation, and (2) complete failure on 60% of low-contrast myoblast samples where texture features proved unreliable. These constraints are particularly problematic for longitudinal studies requiring both speed and consistency across imaging sessions. The error message ”Either no cells found or all cells are touching the border” typically occurs in image analysis or cell segmentation tasks when the algorithm fails to properly identify cells or detects cells that are too close to the image edges.
@@ 3.2 Acquiring Dataset:
The dataset consists of 783 (256x256) bright field microscopy images and 85 phase contrast microscopy images of unstained live cells in culture medium. Therefore, the total number of instances is 868. The dataset is obtained by our research group.
@@ 3.3.2 Data Augmentation
We implemented an extensive augmentation pipeline using the albumentations library (v1.4.3). Each image–mask pair was expanded into 40 variants, yielding a total of 320 training samples. The augmentation strategy consisted of two sequential transformation stages: a primary photometric transformation followed by a secondary geometric transformation.
@@
All geometric transformations were applied using nearest-neighbor interpolation to ensure that the masks remained binary and topologically consistent. Additionally, mask binarization was enforced using a fixed threshold of 127 on 8-bit grayscale images. Each augmented sample was composed of one primary transformation (or left unaltered for baseline copies) followed by one randomly selected secondary transformation. This dual-stage augmentation pipeline was specifically designed to simulate the types of variation commonly observed in brightfield microscopy, including optical artifacts, biological heterogeneity, and staining inconsistencies. As a result, the augmented dataset contributed significantly to the model’s ability to generalize, helping it achieve a test accuracy of 91% without overfitting to the small original training set.
@@ 3.5 Training Strategy:
The training is implemented on $256\times 256\times 3$ image patches with a batch size of 16, optimized for GPU memory, and trained for up to 50 epochs with early stopping (patience=5=5 epochs) to improve convergence and avoid overfitting. Performance is evaluated using pixel-level metrics (accuracy, precision, recall) and segmentation quality measures (Dice coefficient, Jaccard index), with additional error analysis through false/true positive/negative rates. Key advantages include memory-efficient attention mechanisms (MODEL-2 & MODEL-3), adaptive hard mining for dynamic difficulty assessment (MODEL-1 & MODEL-2), boundary-aware loss for edge optimization, and mixed precision (MODEL-2 & MODEL-3) training to enable larger batch sizes without compromising accuracy. Therefore total 6 models (Model-1 , 2, 3 with two loss functions, BCE Dice Loss & Focal Dice Boundary Loss) were trained 45 epochs, and ”.keras” files for each epoch were saved. Finally 14 classifiers were selected for voting based on the Best Validation Metrics & Stable Performance.
@@ 4.2 External Validation on LIVECell Dataset
The segmentation model, trained on a dataset containing only 10% phase-contrast images (85 out of 868 $\times256$ training instances), was evaluated on 3,188 annotated images from the LIVECell dataset [^3] https://sartorius-research.github.io/LIVECell/, excluding 8 corrupted samples. The name of the excluded images are:
@@ 4.4 Compute Resource:
Our model delivers high performance with minimal computational requirements, unlike state-of-the-art (SOTA) approaches such as Mesmer [35] (trained on NVIDIA V100 with 32GB VRAM for  72 hours) and Cellpose [43] (benchmarked on high-end GPUs like RTX 2080 Ti and RTX 3090). While operating efficiently on GPUs with just 8–13.7GB memory, our method significantly outperforms SSL [37] (0.12–0.25 IoU, 58s/image on CPU) and Cellpose-SAM [47] (0.31–0.38 IoU, 12s/image on 16GB GPU), achieving 0.67–0.72 IoU at 3.5s/image. This demonstrates superior accuracy and speed with lower hardware demands, making our model highly practical for resource-constrained environments.
@@ 5 Conclusion
Despite its lightweight design, the patch-based inference may appear slow (3–4s/image for resolution $2048\times 1536$) for real-time segmentation (where benchmark is less than a second) limiting real-time use. Boundary localization remains a weakness, as indicated by high Hausdorff distances in some cases. False positives near boundaries or debris occasionally affect precision, pointing to a conservative segmentation tendency.
Notably, the pipeline achieves over 400% higher F1 scores than CellPose and StarDist on bright-field data, without relying on nuclear stains or large annotated datasets. It performs especially well on A172 cells (F1: 0.95+) while identifying improvement areas for challenging types like SHSY5Y (F1: 0.84).
@@ Acknowledgments
The authors would like to thank ITMO University for supporting this research.
@@
The work was completed within the research topic No. 640103, *Development of methods for automated processing and analysis of optical and atomic force microscopy images using machine learning techniques*.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper [^4] This publication is a preprint and has not been peer-reviewed. It is made available under the selected license. Any unauthorized use or violation of the license terms is strictly prohibited and will be the responsibility of the individual or entity engaging in such use.,.

### v2

@@ Abstract
Notably, the model was trained primarily on bright-field images with limited exposure to phase-contrast microscopy (<<20%), yet it generalized effectively to the phase-contrast LIVECell dataset, demonstrating modality, robustness and strong performance. This highlights its potential for real-world laboratory deployment across imaging conditions.
The model requires minimal compute power and is adaptable using basic deep learning setups such as Google Colab, making it practical for training on other cell variants. Our pipeline outperforms existing methods in robustness and precision for bright-field microscopy segmentation. The code and dataset are available for reproducibility  [^1] This paper is under review. The full dataset and code used in this study may be provided upon reasonable request to the corresponding author.
@@ 1 Introduction
Importantly, while our training set was composed primarily of bright-field images, it included a small proportion $(<20\%)$ of phase-contrast images, allowing us to evaluate the model’s robustness across imaging modalities. The model generalized effectively to phase-contrast images in the LIVECell dataset, demonstrating its adaptability and cross-modality potential.
@@ 2 Related Work
The work by [37] represents a significant shift toward annotation-free segmentation through optical flow-based pseudo-labeling. While achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images, our evaluation revealed critical limitations for bright-field applications: (1) processing times of 50–60 seconds per image due to iterative optical flow computation, and (2) complete failure on 60% of low-contrast myoblast samples where texture features proved unreliable. These constraints are particularly problematic for longitudinal studies requiring both speed and consistency across imaging sessions. The error message “Either no cells found or all cells are touching the border” typically occurs in image analysis or cell segmentation tasks when the algorithm fails to properly identify cells or detects cells that are too close to the image edges.
@@ 3.2 Acquiring Dataset:
The dataset consists of 697 (256x256) bright field microscopy images and 160 phase contrast microscopy images of unstained live cells in culture medium. Therefore, the total number of instances is 857. The dataset is obtained by our research group.
@@ 3.3.2 Data Augmentation
We implemented an extensive augmentation pipeline using the albumentations library (v1.4.3). Each image–mask pair was expanded into 21 variants, yielding a total of 17997 training samples. The augmentation strategy consisted of two sequential transformation stages: a primary photometric transformation followed by a secondary geometric transformation.
@@
All geometric transformations were applied using nearest-neighbor interpolation to ensure that the masks remained binary and topologically consistent. Additionally, mask binarization was enforced using a fixed threshold of 127 on 8-bit grayscale images. Each augmented sample was composed of one primary transformation (or left unaltered for baseline copies) followed by one randomly selected secondary transformation. This dual-stage augmentation pipeline was specifically designed to simulate the types of variation commonly observed in brightfield microscopy, including optical artifacts, biological heterogeneity, and staining inconsistencies. As a result, the augmented dataset contributed significantly to the model’s ability to generalize, helping it achieve a test accuracy of 93% without overfitting to the small original training set.
@@ 3.5 Training Strategy:
The training is implemented on $256\times 256\times 3$ image patches with a batch size of 16, optimized for GPU memory, and trained for up to 45 epochs with early stopping (patience=5=5 epochs) to improve convergence and avoid overfitting. Performance is evaluated using pixel-level metrics (accuracy, precision, recall) and segmentation quality measures (Dice coefficient, Jaccard index), with additional error analysis through false/true positive/negative rates. Key advantages include memory-efficient attention mechanisms (MODEL-2 & MODEL-3), adaptive hard mining for dynamic difficulty assessment (MODEL-1 & MODEL-2), boundary-aware loss for edge optimization, and mixed precision (MODEL-2 & MODEL-3) training to enable larger batch sizes without compromising accuracy. Therefore total 6 models (Model-1 , 2, 3 with two loss functions, BCE Dice Loss & Focal Dice Boundary Loss) were trained 45 epochs, and ”.keras” files for each epoch were saved. Finally 14 classifiers were selected for voting based on the Best Validation Metrics & Stable Performance.
@@ 4.2 External Validation on LIVECell Dataset
The segmentation model, trained on a dataset containing only 20% phase-contrast images ( out of 857; $\times256$ training instances), was evaluated on 3,188 annotated images from the LIVECell dataset [^3] https://sartorius-research.github.io/LIVECell/, excluding 8 corrupted samples. The name of the excluded images are:
@@ 4.4 Compute Resource:
Our model delivers high performance with minimal computational requirements, unlike state-of-the-art (SOTA) approaches such as Mesmer [35] (trained on NVIDIA V100 with 32GB VRAM for  72 hours) and Cellpose [43] (benchmarked on high-end GPUs like RTX 2080 Ti and RTX 3090). While operating efficiently on GPUs with just 8–13.7GB memory, our method significantly outperforms SSL [37] (0 - 0.133 IoU, 50 - 60s/image on CPU) and Cellpose-SAM [47] (0.087 - 0.49 IoU, 24 - 26s/image on 16GB GPU), achieving 0.542–0.705 IoU at 13 - 15s/image for images sized 1536 $\times$ 2048. This demonstrates superior accuracy and speed with lower hardware demands, making our model highly practical for resource-constrained environments.
@@ 5 Conclusion
Despite its lightweight design, the patch-based inference may appear slow (3–4s/image for resolution $704\times 520$) for real-time segmentation (where benchmark is less than a second) limiting real-time use. Boundary localization remains a weakness, as indicated by high Hausdorff distances in some cases. False positives near boundaries or debris occasionally affect precision, pointing to a conservative segmentation tendency.
Notably, the pipeline achieves approximately 429% higher F1 scores than StarDist and 48% higher F1 scores than CellPose-SAM on bright-field data, without relying on nuclear stains or large annotated datasets. It performs especially well on A172 cells (mean F1: 0.95 $\pm$ 0.02) while identifying improvement areas for challenging types like SHSY5Y (mean F1: 0.84 $\pm$ 0.05).
@@ Acknowledgments
The authors would like to thank Svetlana Ulasevich for providing microscopy images and for fruitful discussions of the domain area.
@@
The research was supported by ITMO University Research Projects in AI Initiative (RPAII) (project #640103, Development of methods for automated processing and analysis of optical and atomic force microscopy images using machine learning techniques)*.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.